Operations Report – January 02nd

  

 

Crew 218 Operations Report 2-Jan-20

SOL: 11

Name of person filing report: Luz Maria Agudelo Urrego

Non-nominal systems: NA

Generator: run

Hours run: 12.5

From what time last night: 1900

To what time this morning: 0730

List any additional daytime hours when the generator was run: N/A

Solar— SOC % (Before the generator is run at night: 64%

Diesel Reading – 74%

Station Propane Reading – 62%

Ethanol Free Gasoline: N/A

Water (loft tank) (gal): 36

Water Meter (units): 0147236.5

Water (static tank) (gal): 270

Static to Loft Pump used – No

Water in Green Hab (gal): 80

Water in Science Dome (gal): 0

Toilet tank emptied: No

Deimos rover used: Still in the workshop

Hours:

Beginning charge:

Ending charge:

Currently charging:

Sojourner rover used: Assigned to director

Hours:

Beginning charge:

Ending charge:

Currently charging:

Spirit rover used: EVA 10

Hours: 121.7

Beginning charge: (Before EVA): 100%

Ending charge: (On return from EVA, before recharging): 86%

Currently charging: Yes

Spirit rover used: EVA 11

Hours: 122.2

Beginning charge: (Before EVA): 100%

Ending charge: (On return from EVA, before recharging): 57%

Currently charging: Yes

Opportunity rover used: Still in the workshop

Hours:

Beginning charge:

Ending charge:

Currently charging:

Curiosity rover used: EVA 10

Hours: 127

Beginning charge: 100%

Ending charge: (On return from EVA, before charging): 88%

Currently charging: Yes

Curiosity rover used: EVA 11

Hours: 127.5

Beginning charge: 100%

Ending charge: (On return from EVA, before charging): 61%

Currently charging: Yes

Notes on rovers: Opportunity and Deimos off-site for maintenance.

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): No

Reason for use: N/A

Oil Added? No

# Hours the ATVs were used today: 0

Notes on ATVs: N/A

HabCar used, and why, where? No

CrewCar used, and why, where? No

General notes and comments: Nominal

Summary of the internet: Nominal

Summary of suits and radios: We received instruction from Scott Davis about how to fix Suit N°6 on the two-piece space simulator suits. We will fix the suit tomorrow, January 3rd, and let you know how it goes.

Summary of Hab operations: The pipe from the static tank through the loft tank is still frozen. The crew did not fill out the Loft Tank by carrying water buckets from the Static Tank since we still have water in the Loft Tank.

Summary of Science Dome operations: Nominal

Summary of RAM operations: Nominal

Summary of any observatory issues: Nominal

Summary of health and safety issues: Nominal

Questions, concerns, and requests to Mission Support: NA

Luz Ma
ENG
Crew 218